# Anti-correlation between excitations and locally-favored structures in glass-forming systems

### v1

@@ II.5 Calculation of the separation between LFS and excitations
If an LFS particle fits the criterion of “long-lived”, we label it as “LFS” throughout the whole trajectory. For excitations, we label the particle as “EX” during $[t_{0}-t_{a},t_{0}+t_{a}]$, where $t_{0}$ is when the excitation happens. Then we calculate the conditional probability $P(EX|LFS)$ of each frame, and the probability of finding excitations $P(EX)$ of each frame. Thus, we can use $P(EX|LFS)\ P(EX)$ to indicate the probability of excitations overlapping with LFS. For each temperature, we take the average $P(EX|LFS)/P(EX)$ of all frames in 10 trajectories.
@@
Figure 3: Representative snapshots of high temperature (a) and low temperature (b) in KA2:1. Particles in LFS are rendered in blue and excitations are rendered in red. (c) A representative snapshot of the experimental colloidal system.
Then we compute the average distance of all excitations in all frames in all trajectories. We choose these different time criteria between the LFS and excitations because in the case of an excitation occurring inside in LFS, its movement would naturally distort the LFS and likely lead to the LFS no longer being detected. If the LFS is no longer detected, then its separation from the excitation would not be counted, which would have severe consequences for our analysis. We have found that the method we implemented avoids this problem.
@@ III Results
We begin our results section by discussing the glassforming behavior of the systems considered. We probe the increase in relaxation via an Angell plot, and evaluate the increase in LFS upon supercooling. We also determine the drop in population of excitations with supercooling. We then move on to consider the spatial relationship of LFS and excitations, starting with some snapshots, before presenting our protocol to determine the separation between excitations and LFS. We then investigate local packing of particles in excitations and LFS.
@@ III.1 The proportion of LFS and excitations at deep supercooling
Fig. 1(a) shows a so-called Angell plot of relaxation time $\tau_{\alpha}$ with respect to inverse temperature $1/T$ for the Kob-Andersen binary (KA) mixture and reduced pressure $Z$ for the experimental hard-sphere colloidal system. Here we fit $\tau_{\alpha}$ with the Vogel–Fulcher–Tamman (VFT) equation,
@@
for simulations and experiments, respectively. Here $D$ is a system-dependent constant. We focus on the region at deeper supercooling than the mode-coupling cross-over as indicated by the blue shading. The mode-coupling crossover was taken from Ortlieb et al. (2023) to be $T_{\mathrm{mct}}=0.55\pm 0.09$ and $0.7\pm 0.1$ for the KA 2:1 and 3:1 systems respectively. For the experiments, we took a volume fraction $\phi_{\mathrm{mct}}=0.58$ van Megen et al. (1998) which corresponds to a reduced pressure $Z_{\mathrm{mct}}=23.23$.
In Fig. 1 we see that with the decrease in temperature, in the case of the simulations, and increase in reduced pressure for the experiments, the proportion of LFS grows [Fig. 1(b)], while the proportion of excitations decreases, consistent with previous results Ortlieb et al. (2023), Fig. 1(c).
@@ III.2 The anti-correlation between LFS and excitations
We further calculate the probability of overlap between LFS and excitations, and it turns out that the probability is generally very small (Fig. 2). Representative snapshots also confirm that LFS and excitations are generally not overlapped (Fig. 3a-c) (indicating that there may be a spatial separation in between.)
@@ III.3 The spatial separation between LFS and excitations
To further probe the anti-correlation between LFS and excitations and their spatial separation, we calculate the distance between LFS and excitations. We compute the distance of an excitation particle to its closest LFS center (denoted by $d$). Figures 4(a) and 5(a) represent some typical scenarios where the excitation particle is inside the LFS or outside the LFS.
@@
We first choose particles in excitations and compute the distance $d$, and the probability distribution $P(d)$ is shown in Fig.  4(b). However, since the population of LFS increases with supercooling [Fig. 1(b)], then one expects that the separation of a particle chosen at random and the nearest LFS particle to it should decrease with supercooling, quite the opposite trend of increasing separation between LFS and excitations that our hypothesis would suggest. This effect seems to lead to a decrease in typical separation between LFS and excitation particles.
@@
To compensate for this effect of the decrease in excitation–LFS separation, which we expect to be driven by the increase in LFS particles with supercooling, we proceed as follows. We pick some random particles and also compute this probability distribution $P_{\mathrm{rand}}(d)$ which we show in Fig. 4(c). By normalizing the LFS-excitation separation distribution by that of the random particles – LFS, we can compensate for the effect of the increasing LFS population at deep supercooling. Thus by calculating $P(d)/P_{\mathrm{rand}}(d)$, we obtain Fig. 4(d), which reveals the distribution of distances between LFS centers and excitations compared with a random particle.
@@
The black dotted line in Fig. 4(d) indicates that the probability of finding excitation particles here is the same as random particles. Thus, compared to random particles, fewer excitations happen inside LFS (position 1), and more excitations outside LFS (positions 2 and 3). The dashed line in Fig. 4(d) is the linear fit $P(d)/P_{\mathrm{rand}}(d)$, indicating that there is a trend that excitations are more likely to occur outside LFS. (This trend may be more obvious at low temperatures). This spatial separation is also observed in the KA3:1 mixture [Fig. 4(e)] and experimental colloidal system (Fig. 5). Thus, for the systems considered, there is a tendency towards excitations occurring outside the LFS upon supercooling.
Figure 4: The spatial relationship between excitations and LFS for the computer simulations. (a) Some typical scenarios of particles inside LFS and outside LFS. $d$ is the distance from one particle and its closest LFS center. (b) The probability distribution of the distance $d$ from one excitation to its closest LFS center. (c) The probability distribution of the distance $d$ from one random particle to its closest LFS center. (d) Scaled excitation–LFS center distance. [(b) normalized by (c)]. That is to say, the effects of the increasing LFS population with supercooling have been removed. Dashed lines in (d) are linear fits. Data above the black dotted line, indicates that excitations are more likely to be found at this distance from the LFS center compared to random particles. (e) The scaled excitation–LFS distance for the KA 3:1 system. That is to say, the sane representation as (d).
@@
Figure 5: The spatial relationship between excitations and LFS for the experiments. (a) Some typical scenarios of particles inside LFS and outside LFS. $d$ is the distance from one particle and its closest LFS center. (b) Scaled excitation–LFS center distance. This corresponds to Figs. 4(d,e).
@@ III.4 Voronoi Cell Volumes
To investigate why excitations are more likely to occur outside the LFS) we compute the volume of the Voronoi cell for each particle. We consider A and B particles for the KA2:1 composition n Fig. 6, where we plot the distribution of Voronoi volumes for particles in LFS and also those in excitations. While the distributions overlap, there is a clear trend of LFS particles having a small Voronoi volume, ie being well packed. The opposite trend is seen for excitations whose larger Voronoi volumes indicate poorer packing. In the case of the smaller B particles, the trend is stronger. We show the case for the KA 3:1 composition in the SM. For the experimental data, errors in coordinate tracking hamper this analysis, moreover for a given coordinate, we do not know the diameter of a given particle in this polydisperse system Ivlev, A. and Lowen, H and Morfill, G. E. and Royall, C. P. (2012); Royall et al. (2023).
Figure 6: Volumes of Voronoi cells for particles in LFS and excitations in the KA 2:1 composition. (a) The probability distribution of A particles. (b) The probability distribution of B particles. The vertical black line is the average of all particles. Solids lines are particles in excitations, dashes are particles in LFS.
@@ IV Discussion
We have investigated the spatial relationship between LFS and excitations in simulations and experiments on model glassformers. This work was motivated by the recent observation of a signficant anti-correlation between excitations and LFS Ortlieb et al. (2023). This seems to be stronger than is the case for earlier work which examined the relation between local structure and dynamic heterogeneity at weaker supercooling Charbonneau et al. (2012); Malins et al. (2013a).
Here we found a significant anti-correlation between long-lived LFS and excitations. This is stronger than that observed previously, Ortlieb et al. (2023) due to our use of long-lived LFS and using inherent state coordinates. Some comments are in order here. In a series of papers Speck and co-workers explored the dynamical phase transition of dynamic facilitation theory in a number of model glassformers in both simulations and experiment Speck et al. (2012); Pinchaipat et al. (2017); Turci et al. (2017, 2018); Royall et al. (2020); Campo and Speck (2019). This transition – obtained using short (a few $\tau_{\alpha}$) and small (typically around 100-200 particles) trajectories – exists between a so-called active phase (similar to the normal supercooled liquid) and an inactive phase, whose dynamics are too slow to be measured on the simulation timescale. This transition is driven by the time–averaged population of LFS. Therefore, the anti-correlation between excitations (a feature of the active phase) and long-lived LFS (which are likley related to the time-averaged LFS of the inactive phase) may be taken as consistent with the dynamical phase transition. We therefore expect that long-lived LFS are some measure of local stability in the supercooled liquid.
@@
It has been found that there is a rather weaker anti-correlation between LFS and Co-operatively Re-arranging Regions (CRRs) which are the elementary units of relaxation in RFOT and Adam-Gibbs thermodynamic-based theories Ortlieb et al. (2023). If we imagine that the long-lived LFS are representative of the emerging solid glass, the enhanced anti-correlation of the excitations with respect to the CRRs raises questions about these relaxation mechanisms. It has been suggested that CRRs may be comprised of many excitations Ortlieb et al. (2023). The lifetimes even of the long-lived LFS are very much smaller than the structural relaxation time $\tau_{\alpha}$ and indeed the timescale associated with CRRs, which reaches $1.1\times 10^{5}$ LJ time units for the lowest temperature studied Ortlieb et al. (2023). Such a timescale of course presents great challenges for the frequency of sampling that we use here. Here, we have focussed on the excitations, but the link between excitations and the much longer timescale CRRs, and the LFS remains an intriguing and challenging topic for a future investigation.
@@ V Conclusion
Our work provides a picture of structural relaxation via excitations occurring in regions of the system outside long-lived LFS. We hope this will stimulate further investigations of the relationship between structure and dynamics in the deeply supercooled regime of glassforming systems which is now accessible.
Acknowledgements.
The authors would like to acknowledge Ludovic Berthier, Gilles Tarjus and Thomas Speck for insightful discussions. DL gratetully acknowledges École Normale Supérieure for financial support. CPR would like to acknowledge the Agence National de Recherche for the provision of the grant DiViNew.
@@ Acknowledgements.
Data and code are available upon reasonable request.
@@ VI Supporting Information
Table S1: VFT parameters fitted to the various systems considered.
@@
Table S2: Scaling of the population of particles in excitations.
Figure S1: The spatial separation in KA3:1 system. (a) The probability distribution of the distance $d$ between excitations and their closest LFS centers. (b) The probability distribution of the distance $d$ between random particles and their closest LFS centers. The ratio of the data plotted in (a) and (b) is shown in Fig. 4(e) in the main text.
Figure S2: The probability distribution of LFS lifetime. (a) KA2:1. (b) KA3:1. (c) Colloidal experiments.

### v2

@@ II.5 Calculation of the separation between LFS and excitations
If a particle is found to belong to a long-lived LFS, we label it as “LFS” throughout the whole trajectory. For excitations, we label the particle as “EX” during $[t_{0}-t_{a},t_{0}+t_{a}]$, where $t_{0}$ is the time at which the excitation occurs. We choose these different time criteria between the LFS and excitations because in the case of an excitation occurring inside an LFS, its movement would naturally distort the LFS and likely lead to the LFS no longer being detected. If the LFS is no longer detected, then its separation from the excitation would not be counted, which would have severe consequences for our analysis. We have found that the method we implemented avoids this problem.
From this labelling, we then calculate the conditional probability that a particle is in an excitation given that it belongs to an LFS $P(EX|LFS)$, and the probability of finding excitations $P(EX)$, both computed in each configuration along the trajectory. The ratio $P(EX|LFS)/P(EX)$ thus indicates the probability of excitations overlapping with LFS. For each temperature, we compute the average of that ratio over the 10 independent trajectories, each comprising 1000 configurations.
Figure 3: Representative snapshots of high (a) and low (b) temperature configurations in KA 2:1. Particles in LFS are rendered in blue and excitations are rendered in red. (c) A representative snapshot of the experimental colloidal system.
@@ III Results
We begin our results section by discussing the glass-forming behavior of the systems considered. We probe the increase in relaxation via an Angell plot, and evaluate the increase in LFS upon supercooling. We also determine the drop in fraction of particles in excitations with supercooling. We then move on to consider the spatial relationship of LFS and excitations, starting with some snapshots, before presenting our protocol to determine the separation between excitations and LFS. We then investigate the local packing of particles in excitations and LFS.
@@ III.1 The proportion of LFS and excitations at deep supercooling
Figure 1(a) shows a so-called Angell plot of relaxation time $\tau_{\alpha}$ with respect to inverse temperature $1/T$ for the Kob-Andersen binary (KA) mixture and reduced pressure $Z$ for the experimental hard-sphere colloidal system. Here we fit $\tau_{\alpha}$ with the Vogel–Fulcher–Tamman (VFT) equation,
@@
for simulations and experiments, respectively. Here $D$ is a system-dependent constant. We focus on the regime at deeper supercooling than the mode-coupling crossover as indicated by the blue shading. The mode-coupling crossover is taken from Ortlieb et al. (2023) to be $T_{\mathrm{mct}}=0.55\pm 0.09$ and $0.7\pm 0.1$ for the KA 2:1 and 3:1 systems respectively. For the experiments, we take a volume fraction $\phi_{\mathrm{mct}}=0.58$ van Megen et al. (1998) which corresponds to a reduced pressure $Z_{\mathrm{mct}}=23.23$.
We now turn to the evolution of the population of LFS, defined as the number of long-lived LFS $N_{\textrm{lfs}}$ divided by the number of particles $N$. In Fig. 1(b) we see that the population of LFS grows with supercooling in all systems. At the lowest temperature, one particle in three belongs to a long-lived LFS.
We define the concentration of excitations $c_{a}$ as the fraction of particles identified as excitations during the whole trajectory. Dynamic facilitation theory predicts it to decrease exponentially with the inverse temperature Keys et al. (2011) which was verified at deeper supercooling in Ref. Ortlieb et al. (2023). We show in Fig. 1(c) that the data obtained on the IS trajectories is consistent with such scaling.
Figure 4: Renderings of some typical examples of excitations inside LFS (a) and outside LFS (b,c), with $d$ the distance from an excitation particle (red) and its closest LFS center (cyan).
Figure 5: The spatial relationship between excitations and LFS for the computer simulations. Shown is the probability distribution of the distance $d$ from the LFS center as expressed in Fig. 4. We consider the following cases. (a) Excitation to closest LFS center. (b) A randomly chosen particle to the closest LFS center. (c) The ratio of (a) and (b). That is to say, normalizing the excitation–LFS center distance by a randomly chosen particle–LFS distance, which accounts the increase in LFS population with supercooling. Dashed lines are linear fits. (d) Normalized excitation–LFS distance for the KA 3:1 mixture.
@@ III.2 The anti-correlation between LFS and excitations
We further calculate the probability of overlap between LFS and excitations. It turns out that this probability is generally very small (Fig. 2). Representative snapshots shown in Fig. 3 also confirm that LFS and excitations generally do not overlap, indicating that there may be a spatial separation in between.
@@ III.3 The spatial separation between LFS and excitations
To further probe the anti-correlation between LFS and excitations and their spatial separation, we calculate the distance between LFS and excitations. We compute the distance, noted $d$, of an excitation particle to its closest LFS center. Figures 4 and 6(a) render some typical scenarios where the excitation particle is inside or outside an LFS, for simulations and experiments respectively.
@@
We compute the probability distribution $P_{\mathrm{ex}}(d)$, shown in Fig.  5(a). However, since the population of LFS increases with supercooling [Fig. 1(b)], one expects that the distance between a particle chosen at random and its nearest LFS particle to decrease with supercooling. This is the opposite of our hypothesis of an increasing separation between LFS and excitations. This effect seems to lead to a decrease in typical separation between LFS and excitation particles as shown in Fig. 5(a).
@@
To compensate for this effect of the decrease in excitation–LFS separation, which we expect to be driven by the increase in LFS particles with supercooling, we proceed as follows. We pick some random particles and also compute the probability distribution $P_{\mathrm{rand}}(d)$ of its distance to the closest LFS center, which we show in Fig. 5(b). This plot is related to the pair distribution function, but constrained to one particle being a LFS center. It therefore exhibits peaks at multiples of the particle diameters. However, it is cut-off at large distance because one considers the closest LFS center. This cutoff value decreases with temperature, which is consistent with the increase in LFS density reported in Fig. 1(b).
By normalizing the excitation-LFS separation distribution by that of the random particles-LFS, we can compensate for the effect of the increasing LFS population at deep supercooling. Thus by calculating $P_{\mathrm{ex}}(d)/P_{\mathrm{rand}}(d)$, we obtain Fig. 5(c), which reveals the distribution of distances between LFS centers and excitations compared with a random particle. The horizontal dotted line $P_{\mathrm{ex}}(d)/P_{\mathrm{rand}}(d)=1$ in Fig. 5(c) separates spatial regions where the probability of finding an excitation is smaller or greater than finding a random particle. We see that the probability to have an excitation inside the LFS, corresponding to $d<1$ or Fig. 4(a), is zero. Instead, we observe that excitations are increasingly ‘expelled’ from the center of LFS as temperature decreases. We indeed identify two peaks above the horizontal line, located around $d\sim 1.4$ and $d\sim 2.2$, with the second peak higher and increasing with supercooling, indicating a larger probability to find excitations further away from LFS. The location of these peaks correspond to excitation-LFS distances shown in Fig. 4(b) and (c), respectively.
@@
In order to highlight this increasing spatial separation between excitations and LFS, we fit the ratio of probabilities by a linear function, shown with dashed lines. We find that there is a trend that excitations are more likely to occur outside LFS, and even so at low temperature. This spatial separation between LFS and excitations is also observed in the KA 3:1 mixture [Fig. 5(d)] and experimental colloidal system (Fig. 6). Thus, for all systems considered here, there is a tendency towards excitations occurring increasingly further away from LFS upon supercooling. This is the main result of our study.
@@
Figure 6: The spatial relationship between excitations and LFS for the experiments. (a) Some typical scenarios of particles inside LFS and outside LFS, with $d$ the distance from one particle and its closest LFS center. (b) Scaled excitation–LFS center distance. This corresponds to Figs. 5(c,d).
@@ III.4 Voronoi Cell Volumes
To investigate why excitations are more likely to occur outside LFS, we compute the volume of the Voronoi cell for each particle. We consider A and B particles for the KA 2:1 composition Fig. 7, where we plot the distribution of Voronoi volumes for particles in LFS and also those in excitations. While the distributions overlap, there is a clear trend of LFS particles having a small Voronoi volume, i.e. being well packed. The opposite trend is seen for excitations whose larger Voronoi volumes indicate poorer packing. In the case of the smaller B particles, the trend is stronger. We show the case for the KA 3:1 composition in the Supplementary information, see Fig. S3. For the experimental data, errors in coordinate tracking hamper this analysis, moreover for a given coordinate, we do not know the diameter of a given particle in this polydisperse system Ivlev, A. and Lowen, H and Morfill, G. E. and Royall, C. P. (2012); Royall et al. (2023).
Figure 7: Volumes of Voronoi cells for particles in LFS and excitations in the KA 2:1 composition. The probability distribution of (a) A particles, and (b) B particles. The vertical black line is the average over all (A or B) particles. Solid lines are particles in excitations, dashes are particles in LFS. Data for the KA 3:1 mixture is provided in Fig. S3.
@@ IV Discussion
We have investigated the spatial relationship between LFS and excitations in simulations and experiments on model glassformers. This work was motivated by the recent observation of a significant anti-correlation between excitations and LFS Ortlieb et al. (2023). This seems to be stronger than is the case for earlier work which examined the relation between local structure and dynamic heterogeneity at weaker supercooling Charbonneau et al. (2012); Malins et al. (2013a).
Here we found a significant anti-correlation between long-lived LFS and excitations. This is stronger than that observed previously Ortlieb et al. (2023), due to our use of long-lived LFS and using inherent state coordinates. Some comments are in order here. In a series of papers Speck and co-workers explored the dynamical phase transition of dynamic facilitation theory in a number of model glassformers in both simulations and experiment Speck et al. (2012); Pinchaipat et al. (2017); Turci et al. (2017, 2018); Royall et al. (2020); Campo and Speck (2020). This transition – obtained using short (a few $\tau_{\alpha}$) and small (typically around 100-200 particles) trajectories – exists between a so-called active phase (similar to the normal supercooled liquid) and an inactive phase, whose dynamics are too slow to be measured on the simulation timescale. In the work of Speck and collaborators, this transition is driven by the time–averaged population of LFS, (or by dynamics, as obtained previously Hedges et al. (2009); Speck and Chandler (2012)). Therefore, the anti-correlation between excitations (a feature of the active phase) and long-lived LFS (which are likely related to the time-averaged LFS of the inactive phase) may be taken as consistent with the dynamical phase transition. We therefore expect that long-lived LFS are some measure of local stability in the supercooled liquid.
@@
It has been found that there is a rather weaker anti-correlation between LFS and cooperatively re-arranging regions (CRRs) which are the elementary units of relaxation in RFOT and Adam-Gibbs thermodynamic-based theories Ortlieb et al. (2023). If we imagine that the long-lived LFS are representative of the emerging solid glass, the enhanced anti-correlation of the excitations with respect to the CRRs raises questions about these relaxation mechanisms. It has been suggested that CRRs may be comprised of many excitations Ortlieb et al. (2023). The lifetimes even of the long-lived LFS are very much smaller than the structural relaxation time $\tau_{\alpha}$ and indeed the timescale associated with CRRs, which reaches $1.1\times 10^{5}$ LJ time units for the lowest temperature studied Ortlieb et al. (2023). Such a timescale of course presents great challenges for the frequency of sampling that we use here. Here, we have focused on the excitations, but the link between excitations and the much longer timescale CRRs, and the LFS remains an intriguing and challenging topic for a future investigation.
@@ V Conclusion
Our work provides a picture of structural relaxation via excitations occurring in regions of the system outside long-lived LFS. We hope this will stimulate further investigations of the relationship between structure and dynamics in the deeply supercooled regime of glass-forming systems which is now accessible.
acknowledgments
The authors would like to acknowledge Ludovic Berthier, Thomas Speck and Gilles Tarjus for insightful discussions. DL gratetully acknowledges École Normale Supérieure for financial support. CPR would like to acknowledge the Agence National de Recherche for the provision of the grant DiViNew.
@@ acknowledgments
Data and codes are available upon reasonable request to the corresponding author.
@@ VI Supporting Information
Table 1: VFT parameters fitted to the various systems considered.
@@
Table 2: Scaling of the fraction of particles in excitations.
Figure S1: The probability distribution of LFS lifetime in simulations of the KA system at (a) 2:1, (b) 3:1 composition, and (c) colloidal experiments. LFS with a lifetime greater than the vertical dashed red lines are defined as long-lived.
Figure S2: The spatial separation in the KA 3:1 system. (a) The probability distribution of the distance $d$ between excitations and their closest LFS centers. (b) The probability distribution of the distance $d$ between random particles and their closest LFS centers. The ratio of the data plotted in (a) and (b) is shown in Fig. 5(d) in the main text.
Figure S3: Volumes of Voronoi cells for particles in LFS and excitations in the KA 3:1 composition. (a) The probability distribution of A particles. (b) The probability distribution of B particles. The vertical black line is the average of all particles. Solids lines are particles in excitations, dashes are particles in LFS.